# Explainable and Class-Revealing Signal Feature Extraction via Scattering Transform and Constrained Zeroth-Order Optimization

### v2

@@ II Our Proposed Method
Perhaps, it is a good idea to describe the ST a bit more precisely here. Since we only consider the second-layer ST coefficients, we define it for a given input 1D signal $\boldsymbol{x}\in\mathbb{R}^{d}$ as follows:
@@
Also, we should explain Step 3 above a bit more here. To solve the minimization problem Eq. (1), we use the so-called zeroth-order (ZO) or derivative-free optimization [13, 14, 15]. For our proposed method, it is better to use ZO optimization algorithms than the popular first-order (FO) optimization algorithms that require computing the gradient of the objective function to be minimized. This is because the gradient $\nabla{\mathcal{S}}[\boldsymbol{x}]$ is highly discontinuous, leading update steps to converge slowly, as shown in [8, Chap. 4]. In this paper, we use the popular ZO method called Differential Evolution (DE) [16, 17, 18]. We also note that we perform the updates of solution candidates in the frequency domain (via DCT) with the pink (a.k.a. $1/f$) noise as the initial randomized candidates instead of time domain updates starting from white noise. This is because the updates of solution candidates in the time domain are too local in time and many naturally-occurring signals follow the $1/f$ power spectra [19]; see also [8, Chap. 4]. The DE is certainly not the only ZO method to use for our problems; see more discussion of the choice of ZO methods in Section IV.
@@ III Examples
In this section, we will demonstrate the usefulness of our proposed method using two well-known synthetic signal classification problems.
@@
(a) Five samples/class in the “CBF” classification problem
(b) $\boldsymbol{{\beta}}$ coefficients of “CBF” classes
(c) Solutions of Eq. (1) for the “CBF” classification problem
@@ III-B Triangular Waveform Classification
(a) Five samples/class of the triangular waveform problem
(b) $\boldsymbol{{\beta}}$ coefficients of the three classes
(c) Solutions of Eq. (1) for the triangular waveform problem
@@
This is another well-known classification problem, first appeared in the classic book [28, Sec. 2.6.2]. The dimensionality (or length) of each signal was extended from the original $21$ in [28] to $128$ in order to fully utilize the power of the ST: $21$ was just too short for cascades of wavelet convolutions. Three classes of signals were generated by the following formulas:
@@
where $i=1,\ldots,128$, $h_{1}(i)=\max(6-|i-43|/7,0)$, $h_{2}(i)=h_{1}(i-21)$, $h_{3}(i)=h_{1}(i-42)$, $u$ is a uniform random variable on the interval $(0,1)$, and $\epsilon(i)$ are the standard normal variates. Simply speaking, each class consists of random convex linear combination of two triangles with additive white Gaussian noise. Notice that the noiseless versions of these three classes form an edge of a triangular manifold in $\mathbb{R}^{128}$ where each vertex is those three triangles $\boldsymbol{h}_{k}$, $k=1,2,3$. Hence, the noisy versions are distributed within Gaussian balls around this triangular manifold. Figure 2(a) shows five sample waveforms from each class; Fig. 2(b) shows the sparse $\boldsymbol{{\beta}}_{k}$ coefficient vectors; and Fig. 2(c) displays the extracted feature/the best input pattern for each class; the red, green, and blue triangles in the background indicate the vectors $\boldsymbol{h}_{1}$, $\boldsymbol{h}_{2}$, and $\boldsymbol{h}_{3}$, respectively. We used the same parameter settings as those of the “CBF” signal classification problem. The classification rate using the ST plus MLR was $88.9$% using $1000$ newly generated signals per class as the test dataset while each of the estimated features in Fig. 2(c) gave us essentially perfect classification rate of $99.99$%. In other words, these estimated features essentially discovered the nature of class signal generation mechanism as in the “CBF” signal classification problem. The Class 1 and the Class 3 features in the bottom and top rows in Fig. 2(c) indicate the signal energy concentrates around the apices of $\{\boldsymbol{h}_{1},\boldsymbol{h}_{2}\}$ and $\{\boldsymbol{h}_{2},\boldsymbol{h}_{3}\}$, respectively while the Class 2 feature shown in the middle row of Fig. 2(c) does not have such energy concentration around $\boldsymbol{h}_{2}$, indicating that the signals of this class do not have the contribution from $\boldsymbol{h}_{2}$. This is quite interesting: if the Class 2 feature had energy concentration around $\{\boldsymbol{h}_{1},\boldsymbol{h}_{3}\}$, it would have generated worse classification rate because it would confuse some Class 1 and Class 3 signals; rather our algorithm decided to use the “nonexistence” of the $\boldsymbol{h}_{2}$ component. This has given us a serendipity: in order to understand some competing class signals, it would be better to consider feature “suppressors” instead of feature extractors!
@@ IV Summary and Discussion
In this paper, we described our effort to understand nonlinear signal features computed by ST and the MLR classifier with the Lasso penalty with the help of the ZO optimization with constraints on sparsity and smoothness of the features in the original time domain.. Our numerical results on the two well-known synthetic signal classification problems were quite promising.
We have used the DE algorithm as our ZO method to solve Eq. (1). Clearly, other methods may work well with even faster computational speed than the DE algorithm. Hence, it is important to evaluate various ZO algorithms that is robust and computationally efficient. There are two different strategies in ZO optimization: single-particle methods and multi-particle methods. The advantage of multi-particle methods [29, 30, 31] is its population-based adaptive and evolutionary solution-search capability while the disadvantage is their difficulty to derive theoretical guarantee of convergence. Hence, we will also investigate the single-particle methods, which are theoretically more tenable and can fully utilize the advanced FO optimization algorithms once the gradient estimation is done at each iteration. In particular, we plan to investigate the so-called ZO proximal methods [32, 33] because these can handle the constraints consisting of a nonconvex and smooth function and a nonsmooth and convex function like our Eq. (1). Our preliminary numerical experiments using the PRIMA.jl [34] is quite promising: it generated similar results as those in Section III with faster computational speed than the DE algorithm.
Furthermore, we need to investigate: 1) semi-automatic determination of the optimal values of $\mu$ and $\nu$ in Eq. (1); 2) possible constraints other than sparsity and smoothness—measured by $\ell_{1}$-norm of the signal and $\ell_{2}$-norm of its gradient—are effective; and 3) the influence of the setup of the wavelet filters (e.g., type of wavelets, number of filters, their frequency overlaps, etc.) used in the ST to the estimated class-revealing features.
Finally, we want to note that our optimization strategy Eq. (1) should work in principle with any other learning model as long as it outputs the class probability $p_{k}(\boldsymbol{x})$ for a given input signal $\boldsymbol{x}\in\mathbb{R}^{d}$, e.g., a DL model equipped with “softmax” output units [4, Sec. 6.2.2.3].

### v3

@@ II Our Proposed Method
We now describe the ST a bit more precisely. Since we only consider the second-layer ST coefficients, we define it for a given input 1D signal $\boldsymbol{x}\in\mathbb{R}^{d}$ as follows:
@@
Step 3 needs a bit more explanation here. To solve the minimization problem Eq. (1), we use the so-called zeroth-order (ZO) or derivative-free optimization [13, 14, 15]. For our proposed method, it is better to use ZO optimization algorithms than the popular first-order (FO) optimization algorithms that require computing the gradient of the objective function to be minimized. This is because the gradient $\nabla{\mathcal{S}}[\boldsymbol{x}]$ is highly discontinuous, leading update steps to converge slowly, as shown in [8, Chap. 4]. In this paper, we use the popular ZO method called Differential Evolution (DE) [16, 17, 18]. We also note that we perform the updates of solution candidates in the frequency domain (via DCT) with the pink (a.k.a. $1/f$) noise as the initial randomized candidates instead of time domain updates starting from white noise. This is because the updates of solution candidates in the time domain are too local in time and many naturally-occurring signals follow the $1/f$ power spectra [19]; see also [8, Chap. 4]. The DE is certainly not the only ZO method to use for our problems; see more discussion of the choice of ZO methods in Section IV. Finally, we point out the importance of using the sparsity and the smoothness constraints in Eq. (1): 1) the landscape of Eq. (1) with $\mu=\nu=0$ becomes too rough for the ZO optimization to converge; and 2) the sparsity and the smoothness provide us with easy and intuitive interpretation of signal features.
@@ III Examples
We now demonstrate the usefulness of our proposed method using two synthetic signal classification problems.
@@
(a) Five samples/class in the “CBF” classification problem: “Cylinder” (bottom 5); “Bell” (middle 5); “Funnel” (top 5)
(b) $\boldsymbol{{\beta}}$ coefficients: “Cylinder” (bottom); “Bell” (middle); “Funnel” (top)
(c) Solutions of Eq. (1): “Cylinder” (bottom); “Bell” (middle); “Funnel” (top)
@@ III-B Triangular Waveform Classification
(a) Five samples/class of the triangular waveform problem: Class 1 (bottom 5); Class 2 (middle 5); Class 3 (top 5)
(b) $\boldsymbol{{\beta}}$ coefficients: Class 1 (bottom); Class 2 (middle); Class 3 (top)
(c) Solutions of Eq. (1): Class 1 (bottom); Class 2 (middle); Class 3 (top)
@@
This is another well-known classification problem, first appeared in the classic book [28, Sec. 2.6.2]. The dimensionality (or length) of each signal was extended from the original $21$ in [28] to $128$ in order to fully utilize the power of the ST: $21$ was just too short for cascades of wavelet convolutions. Three classes of signals were generated by:
@@
where $i=1,\ldots,128$, $h_{1}(i)=\max(6-|i-43|/7,0)$, $h_{2}(i)=h_{1}(i-21)$, $h_{3}(i)=h_{1}(i-42)$, $u$ is a uniform random variable on the interval $(0,1)$, and $\epsilon(i)$ are the standard normal variates. Simply speaking, each class consists of random convex linear combination of two triangles with additive white Gaussian noise. Notice that the noiseless version of each class forms an edge of a triangular manifold in $\mathbb{R}^{128}$ whose vertices are those three triangles $\boldsymbol{h}_{k}$, $k=1,2,3$. Hence, the noisy versions are distributed within Gaussian balls around this triangular manifold. Figure 2(a) shows five sample waveforms from each class; Fig. 2(b) shows the sparse $\boldsymbol{{\beta}}_{k}$ coefficient vectors; and Fig. 2(c) displays the extracted feature/the best input pattern for each class; the red, green, and blue triangles in the background indicate the vectors $\boldsymbol{h}_{1}$, $\boldsymbol{h}_{2}$, and $\boldsymbol{h}_{3}$, respectively. We used the same parameter settings as those of the “CBF” signal classification problem. The classification rate using the ST plus MLR was $88.9$% using $1000$ newly generated signals per class as the test dataset while each of the estimated features in Fig. 2(c) gave us essentially perfect classification rate of $99.99$%. In other words, these estimated features essentially discovered the nature of class signal generation mechanism as in the “CBF” signal classification problem. The Class 1 and the Class 3 features in the bottom and top rows in Fig. 2(c) indicate the signal energy concentrates around the apices of $\{\boldsymbol{h}_{1},\boldsymbol{h}_{2}\}$ and $\{\boldsymbol{h}_{2},\boldsymbol{h}_{3}\}$, respectively while the Class 2 feature shown in the middle row of Fig. 2(c) does not have such energy concentration around $\boldsymbol{h}_{2}$, indicating that the signals of this class do not have the contribution from $\boldsymbol{h}_{2}$. This is quite interesting: if the Class 2 feature had energy concentration around $\{\boldsymbol{h}_{1},\boldsymbol{h}_{3}\}$, it would have generated worse classification rate because it would confuse some Class 1 and Class 3 signals; hence our algorithm decided to use the “nonexistence” of the $\boldsymbol{h}_{2}$ component. This has given us a serendipity: in certain situations, it would be better to consider feature “suppressors” instead of feature extractors!
@@ IV Summary and Discussion
We have described our effort to understand nonlinear signal features computed by ST and the MLR classifier using the ZO optimization with constraints on sparsity and smoothness of the features in the original time domain. Our numerical results in Section III were quite promising.
We have used the DE algorithm as our ZO method to solve Eq. (1). Clearly, other methods may work well with even faster computational speed than the DE algorithm. Hence, it is important to evaluate various ZO algorithms that is robust and computationally efficient. There are two different strategies in ZO optimization: single-particle methods and multi-particle methods. The advantage of multi-particle methods [29, 30, 31] is its population-based adaptive and evolutionary solution-search capability while the disadvantage is their difficulty to derive theoretical guarantee of convergence. Hence, we will also investigate the single-particle methods, which are theoretically more tenable and can fully utilize the advanced FO algorithms once the gradient estimation is done at each iteration. In particular, we plan to investigate the so-called ZO proximal methods [32, 33] because these can handle the constraints consisting of a nonconvex and smooth function and a nonsmooth and convex function like our Eq. (1). Our preliminary numerical experiments using the PRIMA.jl [34] is quite promising: it generated similar results as those in Section III with an order of magnitude faster than the DE algorithm.
Furthermore, we will investigate: 1) automatic determination of the optimal values of $\mu$ and $\nu$ in Eq. (1); 2) other effective constraints beyond sparsity and smoothness; and 3) the influence of the wavelet filter parameters (e.g., type of wavelets, number of filters, their frequency overlaps, etc.).
Finally, we note that our optimization strategy Eq. (1) should work in principle with any other learning model as long as it outputs the class probability $p_{k}(\boldsymbol{x})$ for a given input signal $\boldsymbol{x}\in\mathbb{R}^{d}$, e.g., a DL model equipped with “softmax” output units [4, Sec. 6.2.2.3].